# Quantum backreaction in an analogue black hole

### v1

@@ Appendix B Elementary excitations and quantum modes
In a analogue black configuration the flow is upstream subsonic and downstream supersonic. The Doppler effect is thus different in these regions, from which it stems that the asymptotic dispersion relations are also different. In both asymptotic regions the flow is uniform (with density $\rho_{\alpha}$ and velocity $V_{\alpha}$, where $\alpha=u$ in the upstream asymptotic region and $\alpha=d$ downstream, see Appendix A). In these regions the elementary excitations are thus plane waves. Denoting their angular frequency as $\omega$ and their wave vector as $q$ we have [^4] As in Appendix A, throughout this appendix we omit for legibility all the subscripts “GP” and write $V_{\alpha}$ instead of $V_{\scriptscriptstyle\rm GP}^{\alpha}$ for instance.
@@
Figure 3: Upper plot: dispersion relation (95) in the asymptotic upstream subsonic region. Lower plot: dispersion relation in the downstream supersonic region. This plot is shaded to emphasize that this region corresponds to the interior of the analogue black hole. The horizontal dashed line is the angular frequency $\omega$ of a given excitation. The colored points mark the corresponding excitation channels, of wavevectors $q_{0|{\rm out}}(\omega)$, $q_{0|{\rm in}}(\omega)$, $q_{1|{\rm in}}(\omega)$ etc. The arrows indicate the direction of propagation of the different channels. The wave vector $\kappa_{d}$ corresponds to a zero energy channel, see Sec. IV.2.
The propagation channels with note with a label “in” propagate towards the horizon, the ones with label “out” propagate away fro the horizon.
Several propagation modes correspond to the channels identified in Fig. 3. For instance, a mode with we denote as “ingoing” and associate to a quantum operator $\hat{b}_{0}(\omega)$ corresponds to a wave incident in channel $0|{\rm in}$ at energy $\hbar\omega$, scattered onto the horizon to the exit channels $0|{\rm out}$, $1|{\rm out}$ and $2|{\rm out}$ with amplitudes $S_{0,0}(\omega)$, $S_{0,1}(\omega)$ and $S_{0,2}(\omega)$, respectively. There are two other ingoing modes which are associated to operators we denote as $\hat{b}_{1}(\omega)$ and $\hat{b}_{2}(\omega)$. They correspond to modes initiated by a wave incident in channel $1|{\rm in}$ and $2|{\rm in}$, respectively. There are also three outgoing modes associated to scattering processes resulting in the emission of a single wave along one of the three “out” channels $0|{\rm out}$, $1|{\rm out}$, and $2|{\rm out}$. We denote the corresponding quantum operators as $\hat{c}_{0}(\omega)$, $\hat{c}_{1}(\omega)$ and $\hat{c}_{2}(\omega)$. In résumé, the modes, be they ingoing or outgoing, are identified by an index $L\in\{0,1,2\}$ and the channels by an index $\ell\in\{0|{\rm in},0|{\rm out},1|{\rm in},1|{\rm out},2|{\rm in},2|{\rm out}\}$. We also note that the outgoing modes $\hat{c}_{0}$ and $\hat{c}_{2}$ are analogous to the modes denoted in the General Relativity context as the Hawking and partner modes, respectively.
@@
where for legibility we omit the $\omega$ dependence of all the terms. The appearance of annihilation operators $\hat{b}_{2}^{\dagger}$ and $\hat{c}_{2}^{\dagger}$ in (97) reflects the fact that the modes $L=2$ have a negative norm and should accordingly be quantized inverting the usual role of the creation and annihilation operators in order that the elementary excitations satisfy the standard Bose commutation relations [55].
The $3\times 3$ scattering matrix $S(\omega)$ defined in (97) obeys a skew-unitarity relation [63] :
@@ Appendix C Asymptotic source terms in a 1D black hole
where [^5] As in Appendices A and B, throughout this appendix we omit for legibility all the subscripts “GP” and write $V_{\alpha}$ instead of $V_{\scriptscriptstyle\rm GP}^{\alpha}$ for instance. $k_{\alpha}=k_{u}=mV_{u}/\hbar$ is $x<0$ and $k_{\alpha}=k_{d}=mV_{d}/\hbar$ if $x>0$. In this expression the $u_{\scriptscriptstyle L}(x,\omega)$’s and $v_{\scriptscriptstyle L}(x,\omega)$’s are linear combinations of the usual Bogoliubov coefficients involving the coefficient of the scattering matrix (97), see Eqs, (C) and (C) below.
At zero temperature the different contributions to the source terms which appear in Eqs. (68) and (69) without derivative read [64]
$g^{(2)}(x)=$ $\dfrac{2}{\rho_{\alpha}|\phi_{\alpha}|^{4}}\int_{0}^{\infty}\dfrac{d\omega}{2\pi}\sum_{L=0}^{1}\Big{[}|\tilde{v}_{L}|^{2}+{\it Re}(\tilde{u}_{\scriptscriptstyle L}\tilde{v}_{\scriptscriptstyle L}^{*})\Big{]}$ (100)
$\,+\dfrac{2}{\rho_{\alpha}|\phi_{\alpha}|^{4}}\int_{0}^{\Omega}\frac{d\omega}{2\pi}\Big{[}|\tilde{u}_{2}|^{2}+{\it Re}(\tilde{u}_{2}\tilde{v}_{2}^{*})\Big{]},$
@@
The Bogoliubov coefficients $\tilde{\cal U}_{\ell}(\omega)$ and $\tilde{\cal V}_{\ell}(\omega)$ are real; their explicit expression is given in Ref. [54]. Note that the $q_{\ell}$’s and $S_{i,j}$’s in Eqs. (100) and (101) all depend on $\omega$, see Appendix B.
@@
We recall that $q_{0|{\rm out}}(\omega)<0$, whereas $q_{1|{\rm out}}(\omega)$ and $q_{2|{\rm out}}(\omega)$ are both positive, see Fig. 3. Hence $j_{u}<0$ whereas the integrand of $j_{d}$ contains two contributions, one positive, one negative. In practice we find numerically that the second is dominant and that $j_{d}$ is negative. Accordingly to Eqs. (109) and (110) the dimensionless quantities ${\cal J}^{(H)}_{\alpha}$ defined in Eq. (78b) read:
@@
We determine numerically the four quantities ${\cal G}_{\alpha}^{(H)}$ and ${\cal J}_{\alpha}^{(H)}$ ($\alpha=u$ and $d$) from expression (106), (107), (111) and (112) after a numerical computation of the elements of the $S$-matrix [54]. It may be shown that these quantities cancel in the limit $M_{d}\to 1$. This should be expected since in this limit the negative norm channels $2|{\rm out}$ and $2|{\rm in}$ vanish, resulting in a disappearance of Hawking radiation. Also, since in this limit the upper integration point in integrals (106), (107), (108) and (111) tend to zero, it is easy to show that
@@ Appendix E Asymptotic backreaction for $\delta$-peak and flat profile configurations
In the main text we present the modification of the asymptotic Mach numbers for a waterfall configuration. The reason why we put the emphasize on this configuration is that this is the only one which has been realized experimentally so far [3, 61]. In the present appendix we consider two other configurations (dubbed $\delta$-peak and flat profile in Appendix A) in order to check to what extent the waterfall configuration is typical.
@@
In Fig. 4 we represent the modification of the upstream and downstream Mach number due to quantum backreaction in a $\delta$-peak configuration. The gross features of the results are the same as those obtained for the waterfall configuration and presented in Fig. 1. A noticeable difference between the two figures is that the part $\delta M_{d}^{(H)}$ of the downstream modifications caused bu Hawking radiation, while being roughly of the same order in both settings, are positive for the waterfall configuration and negative for the $\delta$-peak configuration (compare the lower panels of Figs. 1 and 4).

### v2

@@ Appendix B Elementary excitations and quantum modes
In a analog black configuration the flow is upstream subsonic and downstream supersonic. The Doppler effect is thus different in these regions, from which it stems that the asymptotic dispersion relations are also different. In both asymptotic regions the flow is uniform (with density $\rho_{\alpha}$ and velocity $V_{\alpha}$, where $\alpha=u$ in the upstream asymptotic region and $\alpha=d$ downstream, see Appendix A). In these regions the elementary excitations are thus plane waves. Denoting their angular frequency as $\omega$ and their wave vector as $q$ we have [^4] As in Appendix A, throughout this appendix we omit for legibility all the subscripts “GP” and write $V_{\alpha}$ instead of $V_{\scriptscriptstyle\rm GP}^{\alpha}$ for instance.
@@
Figure 3: Upper plot: dispersion relation (96) in the asymptotic upstream subsonic region. Lower plot: dispersion relation in the downstream supersonic region. This plot is shaded to emphasize that this region corresponds to the interior of the analog black hole. The horizontal dashed line is the angular frequency $\omega$ of a given excitation. The colored points mark the corresponding excitation channels, of wavevectors $q_{0|{\rm out}}(\omega)$, $q_{0|{\rm in}}(\omega)$, $q_{1|{\rm in}}(\omega)$ etc. The arrows indicate the direction of propagation of the different channels. The wave vector $\kappa_{d}$ corresponds to a zero energy channel, see Sec. IV.2.
The propagation channels with note with a label “in” propagate towards the horizon, the ones with label “out” propagate away from the horizon.
Several propagation modes correspond to the channels identified in Fig. 3. For instance, a mode with we denote as “ingoing” and associate to a quantum operator $\hat{b}_{0}(\omega)$ corresponds to a wave incident in channel $0|{\rm in}$ at energy $\hbar\omega$, scattered onto the horizon to the exit channels $0|{\rm out}$, $1|{\rm out}$ and $2|{\rm out}$ with amplitudes $S_{0,0}(\omega)$, $S_{0,1}(\omega)$ and $S_{0,2}(\omega)$, respectively. There are two other ingoing modes which are associated to operators we denote as $\hat{b}_{1}(\omega)$ and $\hat{b}_{2}(\omega)$. They correspond to modes initiated by a wave incident in channel $1|{\rm in}$ and $2|{\rm in}$, respectively. There are also three outgoing modes associated to scattering processes resulting in the emission of a single wave along one of the three “out” channels $0|{\rm out}$, $1|{\rm out}$, and $2|{\rm out}$. We denote the corresponding quantum operators as $\hat{c}_{0}(\omega)$, $\hat{c}_{1}(\omega)$ and $\hat{c}_{2}(\omega)$. In short, the modes, be they ingoing or outgoing, are identified by an index $L\in\{0,1,2\}$ and the channels by an index $\ell\in\{0|{\rm in},0|{\rm out},1|{\rm in},1|{\rm out},2|{\rm in},2|{\rm out}\}$. We also note that the outgoing modes $\hat{c}_{0}$ and $\hat{c}_{2}$ are analogous to the modes denoted in the General Relativity context as the Hawking and partner modes, respectively.
@@
where for legibility we omit the $\omega$ dependence of all the terms. The appearance of annihilation operators $\hat{b}_{2}^{\dagger}$ and $\hat{c}_{2}^{\dagger}$ in (98) reflects the fact that the modes $L=2$ have a negative norm and should accordingly be quantized inverting the usual role of the creation and annihilation operators in order that the elementary excitations satisfy the standard Bose commutation relations [Blaizot1986].
The $3\times 3$ scattering matrix $S(\omega)$ defined in (98) obeys a skew-unitarity relation [Recati2009] :
@@ Appendix C Asymptotic source terms in a 1D black hole
where [^5] As in Appendixes A and B, throughout this appendix we omit for legibility all the subscripts “GP” and write $V_{\alpha}$ instead of $V_{\scriptscriptstyle\rm GP}^{\alpha}$ for instance. $k_{\alpha}=k_{u}=mV_{u}/\hbar$ is $x<0$ and $k_{\alpha}=k_{d}=mV_{d}/\hbar$ if $x>0$. In this expression the $u_{\scriptscriptstyle L}(x,\omega)$’s and $v_{\scriptscriptstyle L}(x,\omega)$’s are linear combinations of the usual Bogoliubov coefficients involving the coefficient of the scattering matrix (98), see Eqs, (C) and (C) below.
At zero temperature the different contributions to the source terms which appear in Eqs. (69) and (70) without derivative read [Ciliberto2024_PhD]
$g^{(2)}(x)=$ $\dfrac{2}{\rho_{\alpha}|\phi_{\alpha}|^{4}}\int_{0}^{\infty}\dfrac{d\omega}{2\pi}\sum_{L=0}^{1}\Big[|\tilde{v}_{L}|^{2}+\Re(\tilde{u}_{\scriptscriptstyle L}\tilde{v}_{\scriptscriptstyle L}^{*})\Big]$ (101)
$\,+\dfrac{2}{\rho_{\alpha}|\phi_{\alpha}|^{4}}\int_{0}^{\Omega}\frac{d\omega}{2\pi}\Big[|\tilde{u}_{2}|^{2}+\Re(\tilde{u}_{2}\tilde{v}_{2}^{*})\Big],$
@@
The Bogoliubov coefficients $\tilde{\cal U}_{\ell}(\omega)$ and $\tilde{\cal V}_{\ell}(\omega)$ are real; their explicit expression is given in Ref. [Larre2012]. Note that the $q_{\ell}$’s and $S_{i,j}$’s in Eqs. (101) and (102) all depend on $\omega$, see Appendix B.
@@
We recall that $q_{0|{\rm out}}(\omega)<0$, whereas $q_{1|{\rm out}}(\omega)$ and $q_{2|{\rm out}}(\omega)$ are both positive, see Fig. 3. Hence $j_{u}<0$ whereas the integrand of $j_{d}$ contains two contributions, one positive and one negative. In practice we find numerically that the second is dominant and that $j_{d}$ is negative. Accordingly to Eqs. (110) and (111) the dimensionless quantities ${\cal J}^{(H)}_{\alpha}$ defined in Eq. (79b) read:
@@
We determine numerically the four quantities ${\cal G}_{\alpha}^{(H)}$ and ${\cal J}_{\alpha}^{(H)}$ ($\alpha=u$ and $d$) from expression (107), (108), (112) and (113) after a numerical computation of the elements of the $S$-matrix [Larre2012]. It may be shown that these quantities cancel in the limit $M_{d}\to 1$. This should be expected since in this limit the negative norm channels $2|{\rm out}$ and $2|{\rm in}$ vanish, resulting in a disappearance of Hawking radiation. Also, since in this limit the upper integration point in integrals (107), (108), (109) and (112) tend to zero, it is easy to show that
@@ Appendix E Asymptotic backreaction for $\delta$-peak and flat profile configurations
In the main text we present the modification of the asymptotic Mach numbers for a waterfall configuration. The reason why we put the emphasis on this configuration is that this is the only one which has been realized experimentally so far [deNova2019, Kolobov2021]. In this appendix we consider two other configurations (dubbed $\delta$-peak and flat profile in Appendix A) in order to check to what extent the waterfall configuration is typical.
@@
In Fig. 4 we represent the modification of the upstream and downstream Mach numbers due to quantum backreaction in a $\delta$-peak configuration. The gross features of the results are the same as those obtained for the waterfall configuration and presented in Fig. 1. A noticeable difference between the two figures is that the part $\delta M_{d}^{(H)}$ of the downstream modifications caused by Hawking radiation, while being roughly of the same order in both settings, are positive for the waterfall configuration and negative for the $\delta$-peak configuration (compare the lower panels of Figs. 1 and 4).